# Simulation-Based Inference for Probabilistic Galaxy Detection and Deblending

### v1

@@ 1 Introduction
The distribution of matter in the universe reflects the underlying cosmology that governs its evolution. We can probe the matter distribution through cosmic shear, in which foreground matter coherently distorts background galaxies due to weak gravitational lensing. Cosmic shear is a powerful probe that can be used to constrain cosmology (huterer2002weak; kilbinger2015cosmology) and has been applied successfully in cosmological surveys, including the Sloan Digital Sky Survey (SDSS; huff2014shear), Dark Energy Survey (DES; secco2022desy3_shear; amon2022_desy3_shear_calibration; desy6_shear), Subaru Hyper Suprime-Cam (HSC; dalal2023hsc_shear_spectrum; li2023hsc_shear_2pcf), and Kilo-Degree Survey (KiDS; asgari2021kids). Stage-IV dark energy surveys, such as the Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST; lsst2009book; lsst2019overview), will observe an unprecedented number density of sources because of their greater depth than earlier wide-field surveys. State-of-the-art surveys such as LSST, through analyses of cosmological probes such as cosmic shear, will enable unprecedented precision and accuracy in estimates of cosmological parameters. However, cosmological constraints remain limited by systematics arising from astrophysical processes and measurement procedures (weinberg2013cosmic; mandelbaum2018weak).
@@
In cosmological analyses based on galaxy catalogs derived from optical surveys, a leading systematic is blending (melchior2021challenge). Blending refers to the visual overlap of light sources in astronomical images, which biases and degrades galaxy detection and measurement of their properties. For example, blending can impact the distribution of measured galaxy shapes (hoekstra2017study) and photometry (song2018photometry), the selection function of sources (hartlap2011bias), and the estimated redshift distribution of galaxies $n(z)$ (maccrann2022desy3).
@@
Blending already represents a significant source of systematic error in weak lensing measurements in current ground-based Stage-III surveys (e.g., maccrann2022desy3), and it is expected to become more dominant in Stage-IV surveys, given their greater depth and smaller statistical uncertainties. Simulations suggest that between $62\%$ of galaxies in the LSST will be blended to some degree (sanchez2021effects) and between $15\%$ and $30\%$ of galaxies will be unrecognized blends, in which two or more light sources are identified as one (dawson2016ambiguous; troxel2023joint). Unrecognized blends can significantly affect weak-lensing and cosmological measurements. euclid2019undetected finds that introducing a distribution of faint galaxies in a simulation of Euclid VIS images creates a multiplicative shear bias of at least a few times $10^{-3}$. nourbaksh2021effects uses simulations to forecast the effect of unrecognized blends on the derived structured growth parameter $S_{8}$ for LSST. They find that blending introduces a bias greater than the $2\sigma$ statistical error in estimates of this parameter.
@@
One of the most widely used tools for identifying light profiles in an image is SourceExtractor (SExtractor) (bertin1996sextractor), which uses a model-independent image thresholding technique to detect objects (lutz1980algorithm). It can also perform basic segmentation based on a multiple isophotal analysis technique (based on beard1990cosmos). One major limitation of SExtractor is that each pixel can belong to only one object. The SDSS Photo pipeline (lupton2005sdss) expands upon SourceExtractor, allows for overlap between objects, and estimates the portion of flux in a pixel belonging to each object. Neither SourceExtractor nor the SDSS pipeline can account for multiband information.
@@
Deblenders developed during the last decade, such as MuSCADeT (joseph2016multi) and its successor Scarlet (melchior2018scarlet), build joint nonparametric models of multiband image data. Here, the light sources are not specified using a parametric model, such as a Sérsic fit, but rather by an algorithm that enforces constraints on the light distribution. In particular, Scarlet can impose an arbitrary number of constraints on each source (e.g., symmetry and monotonicity with respect to a source’s centroid), enabling modeling of aspects such as distinct stellar populations in galaxies. More recently, machine-learning-based deblenders have demonstrated high reconstruction accuracy on simulated datasets (arcelin2021deblending; biswas2024madness; sampson2024score; patel2025npe).
@@
Most deblenders are deterministic and do not account for the inherent uncertainty in the properties of a blended source. To overcome this limitation, we introduce a new probabilistic deblender: the Bayesian Light Source Separator (BLISS). BLISS builds on recent ideas in deep generative modeling and variational inference (blei2017variational; kingma2019introduction) to quantify blending uncertainty. Propagating this uncertainty to downstream cosmological analyses has the potential to mitigate blending systematics.
@@
BLISS uses simulation-based inference (SBI) to approximate a posterior distribution of deblended galaxy properties (see Section 3). [^1] As explained in Section 3, the galaxy deblending component of BLISS is deterministic and uses an autoencoder architecture (Section 3.4). Despite this, BLISS can produce posterior distributions of galaxy properties that capture detection uncertainty using the procedure described in Section 4.6. BLISS extends the approach from StarNet (liu2023variational), which uses SBI to detect and measure stars in crowded starfields, to process images containing both stars and galaxies. Algorithmically, inference involves three encoders and a decoder, as illustrated in Figure 1. After training these models, inference can be performed rapidly on a GPU, in just seconds for megapixel images. In this work, we train and evaluate BLISS on simulated single-band images that are representative of typical year-10 LSST wide-field coadd images used for cosmological analyses.
@@ 2 Simulated Images
We designed our dataset of simulated images to mimic year-10 LSST single-band coadds. The galaxies and stars in our image simulations are modeled using parametric light profiles. The source density is taken from previous LSST simulation work (connolly2014catsim; abolfathi2021dc2), and we do not model clustering. All our images are simulated using the GalSim Python package (galsim2015). [^2] https://github.com/GalSim-developers/GalSim
@@ 2.1 Survey Settings
Our studies target a single band, the $i$-band, at the LSST survey pixel scale of 0.2 arcseconds. We use the same parametric point spread function (PSF) model $\pi$ for all our galaxy images, which consists of an atmospheric Kolmogorov component and an optical obscured Airy pattern. We use a FWHM of $0.79$ for the atmospheric component, corresponding to the median zenith seeing for this filter (ivezic2019lsst). We add Gaussian noise to our images, assuming the background-dominated limit, so that their SNR resembles that of LSST year-10 images produced by coaddition. The added noise corresponds to a sky level $b$, which is chosen to be the average expected sky level for LSST year 10. However, the sky level is not added directly to the images; all simulated images are background-subtracted. Our setup closely resembles the simulations used in previous blending work focused on LSST (sanchez2021effects). The exact survey parameters are sourced from the LSST module of SurveyCodex. [^3] SurveyCodex is a Python library designed to archive the parameters of various surveys with references. Many of its parameters were obtained from the WeakLensingDeblending package (sanchez2021wld), in some cases with corrections. This library can be found in the following URL: https://github.com/LSSTDESC/surveycodex
@@ 2.2 Galaxy Properties
The galaxy properties for our image simulations come from a galaxy catalog produced with CatSim, an early version of an end-to-end LSST simulation framework (connolly2014catsim). The galaxies produced by this framework cover the redshift range of $0<z<6$. They are simulated using a parametric model for the light profile of galaxies consisting of bulge and disk components. [^4] The disk component corresponds to a Sérsic profile with index $n=1$ (also known as an exponential profile), and the bulge component corresponds to a Sérsic profile with index $n=4$ (also known as a “de Vaucouleurs” profile) (Sersic1963influence). The model parameters are the total flux, the relative flux of each component, and the sizes and shapes of each component. The two components share the same centroid and orientation. While CatSim also includes an AGN component, we omit it to simplify the modeling of galaxy morphology. The fluxes, shapes, and sizes of galaxies follow a realistic distribution designed to match LSST’s key observables, and are informed by previous SDSS observations as well as galaxy formation literature (bruzual2003stellar; de2006formation; gonzalez2009testing).
For our study, we used the same one-square-degree subset of the CatSim catalog as in sanchez2021effects. This subset of galaxies comprises approximately 850850k galaxies with a limiting $r$-band magnitude of $28$. For our datasets, we impose a magnitude cut in the $i$-band of $27$, which corresponds to galaxies with a median SNR of approximately $3$. This yields approximately 576576k galaxies, which we split into three equal sets for training, validation, and testing.
@@ 2.3 Star Properties
Our datasets include stars to evaluate our deblender in the context of star-galaxy blends. The star distribution and fluxes are derived from the LSST DESC Data Challenge Two (DC2) simulation catalogs (abolfathi2021dc2). Specifically, we use the same star catalog as in previous work on cosmic shear measurements in LSST-like simulations (sheldon2023metadetection). We set our star density to the average over simulated fields, while rejecting fields with density higher than $100\,\text{arcmin}^{-2}$. This results in a star density of $15.5\,\text{arcmin}^{-2}$. We apply a lower limit magnitude cut of $20$ and an upper limit magnitude cut of $27$ in the $i$-band, resulting in approximately 1717k stars. The former cut excludes stars with very high SNRs of over $1000$. We model stars as a point source convolved by the same PSF $\pi$ used by galaxies. We do not simulate saturation or other kinds of image artifacts.
@@ 3 Method
BLISS estimates the posterior distribution of light source properties from a given astronomical image, following the procedure laid out in Figure 1. We start by decomposing the inference problem into subproblems involving overlapping image subregions, which we call “padded tiles” (Section 3.1). These padded tiles are passed into a set of three neural network encoders that are trained via variational inference (ambrogioni2019forward) to return the parameters of distributions characterizing the light source properties in each tile (Section 3.2). First, the detection encoder returns posterior distributions on the number of objects per tile and the positions of objects’ centroids. Sampled centroids from these distributions are used to center the padded tiles (within a pixel, via cropping). Next, the classification encoder takes these centered padded tiles and returns a probability characterizing the type of each source (star or galaxy) (Section 3.3). Finally, the deblending encoder uses these centered padded tiles to return (deterministic) latent galaxy representations that can be used to obtain noiseless reconstructions of deblended galaxy images (Section 3.4). Once the three encoders have been trained, BLISS can process astronomical images of any size with the same PSF and resolution as the training images.
@@ 3.1 Tiling
BLISS, by construction, outputs a prediction for at most one source within a given tile. As a result, if the centroids of two or more sources are present in a single tile, BLISS will output an incomplete catalog for that tile. In particular, if at least one source in the tile is detectable, BLISS will output a set of properties corresponding to a single “hybrid” source. [^8] At training time, BLISS always targets the properties of the brightest source present in a tile. We partially mitigate this by using small tiles of size 5 pixels, which reduces the probability that more than one source centroid lies in any given tile. Given a source density of $175.5\,\rm{arcmin}^{-2}$, this probability is approximately $0.0012$. To be concrete, we expect that approximately $10\%$ of images in our dataset-blend dataset will have at least one tile with two source centroids. In the majority of these occurrences, however, one of the sources involved will be below the detection threshold. [^9] Realistic levels of clustering might significantly impact these assumptions. Clustering increases the number of close pairs of objects at a given number density (euclid2019undetected; maccrann2022desy3). This would increase the proportion of objects that fall in the same tile, thus degrading BLISS performance in the manner mentioned above.
@@ 3.2 Variational inference
BLISS adopts a standard Bayesian model of astronomical images, whereby light source properties are latent random variables, and pixel intensities are observed random variables (cf. brewer2013probabilistic; portillo2017improved; regier2019approximate; buchanan2023markov; stone2023astrophot). Appendix A provides details of our specific Bayesian model.
@@
In our approximation, the inferred tile catalog $\mathcal{Z}_{t}$ contains the light source properties $\{n_{t},\bm{\ell}_{t},b_{t},z_{t}\}$, where $n_{t}\in\{0,1\}$ is the number of sources whose centroids lie in the tile, $\bm{\ell}_{t}\in[0,1]^{2}$ is the centroid position of the tile’s source in units of tile length, $b_{t}\in\{0,1\}$ is the classification of the source as a galaxy ($b_{t}=1$) or star ($b_{t}=0$), and $\bm{z}_{t}\in\mathbb{R}^{8}$ is a latent vector characterizing galaxy morphology and brightness (for more details, see Section 3.4).
@@
are the output of neural networks with weights $\bm{\phi}$ applied to a given padded tile $\mathcal{X}_{t}$. These weights are the variational parameters. We model locations using a bivariate Gaussian in the variational distribution. Samples from this distribution that lie beyond the bounds of a given tile are considered non-detections. [^10] When the true centroid is close to the tile boundary or highly uncertain, there is a significant probability that the posterior samples fall outside the tile. We choose to discard these samples instead of assigning this source to a separate tile due to our current simplifying assumption that tiles are completely independent. Future work will explore sharing information between neighboring tiles to improve the variational approximation. Additionally, we chose the variational distribution for the latent galaxy representation $z_{t}$ to be a delta function. This implies that we only return point estimates for this quantity with no associated uncertainty.
Critically, our method is amortized (kingma2013auto; rezende2014stochastic): the variational parameters $\bm{\phi}$ are shared among tiles (i.e., $\bm{\phi}$ is not indexed by $t$); a single set of neural networks with weights $\bm{\phi}$ is used for any padded tile of any image. If our method were not amortized, as in more classical variational inference approaches (blei2017variational), we would need to run an iterative optimization procedure for each tile, which would make our procedure highly inefficient. Amortization allows us to train a single set of encoders only once, which greatly speeds up inference for large datasets.

### v2

@@ 1 Introduction
The distribution of matter in the universe reflects the underlying cosmology that governs its evolution. We can probe the matter distribution through cosmic shear, in which foreground matter coherently distorts background galaxies due to weak gravitational lensing. Cosmic shear is a powerful probe that can be used to constrain cosmology (Huterer, 2002; Kilbinger, 2015) and has been applied successfully in cosmological surveys, including the Sloan Digital Sky Survey (SDSS; Huff et al. 2014), Dark Energy Survey (DES; Secco et al. 2022; Amon et al. 2022; Yamamoto et al. 2025), Subaru Hyper Suprime-Cam (HSC; Dalal et al. 2023; Li et al. 2023), and Kilo-Degree Survey (KiDS; Asgari et al. 2021). Stage-IV dark energy surveys, such as the Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST; LSST Science Collaboration et al. 2009; Ivezić et al. 2019b), will observe an unprecedented number density of sources because of their greater depth than earlier wide-field surveys. State-of-the-art surveys such as LSST, through analyses of cosmological probes such as cosmic shear, will enable unprecedented precision and accuracy in estimates of cosmological parameters. However, cosmological constraints remain limited by systematics arising from astrophysical processes and measurement procedures (Weinberg et al., 2013; Mandelbaum, 2018).
@@
In cosmological analyses based on galaxy catalogs derived from optical surveys, a leading systematic is blending (Melchior et al., 2021). Blending refers to the visual overlap of light sources in astronomical images, which biases and degrades galaxy detection and measurement of their properties. For example, blending can impact the distribution of measured galaxy shapes (Hoekstra et al., 2017) and photometry (Huang et al., 2018), the selection function of sources (Hartlap et al., 2011), and the estimated redshift distribution of galaxies $n(z)$ (MacCrann et al., 2022).
@@
Blending already represents a significant source of systematic error in weak lensing measurements in current ground-based Stage-III surveys (e.g., MacCrann et al., 2022), and it is expected to become more dominant in Stage-IV surveys, given their greater depth and smaller statistical uncertainties. Simulations suggest that between $62\%$ of galaxies in the LSST will be blended to some degree (Sanchez et al., 2021a) and between $15\%$ and $30\%$ of galaxies will be unrecognized blends, in which two or more light sources are identified as one (Dawson et al., 2015; Troxel et al., 2023). Unrecognized blends can significantly affect weak-lensing and cosmological measurements. Euclid Collaboration et al. (2019) finds that introducing a distribution of faint galaxies in a simulation of Euclid VIS images creates a multiplicative shear bias of at least a few times $10^{-3}$. Nourbakhsh et al. (2021) uses simulations to forecast the effect of unrecognized blends on the derived structured growth parameter $S_{8}$ for LSST. They find that blending introduces a bias greater than the $2\sigma$ statistical error in estimates of this parameter.
@@
One of the most widely used tools for identifying light profiles in an image is SourceExtractor (SExtractor) (Bertin and Arnouts, 1996), which uses a model-independent image thresholding technique to detect objects (Lutz, 1980). It can also perform basic segmentation based on a multiple isophotal analysis technique (based on Beard et al., 1990). One major limitation of SExtractor is that each pixel can belong to only one object. The SDSS Photo pipeline (Lupton et al., 2005) expands upon SourceExtractor, allows for overlap between objects, and estimates the portion of flux in a pixel belonging to each object. Neither SourceExtractor nor the SDSS pipeline can account for multiband information.
@@
Deblenders developed during the last decade, such as MuSCADeT (Joseph et al., 2016) and its successor Scarlet (Melchior et al., 2018), build joint nonparametric models of multiband image data. Here, the light sources are not specified using a parametric model, such as a Sérsic fit, but rather by an algorithm that enforces constraints on the light distribution. In particular, Scarlet can impose an arbitrary number of constraints on each source (e.g., symmetry and monotonicity with respect to a source’s centroid), enabling modeling of aspects such as distinct stellar populations in galaxies. More recently, machine-learning-based deblenders have demonstrated high reconstruction accuracy on simulated datasets (Arcelin et al., 2021; Biswas et al., 2024; Sampson et al., 2024; Patel et al., 2025).
@@
Most deblenders are deterministic and do not account for the inherent uncertainty in the properties of a blended source. To overcome this limitation, we introduce a new probabilistic deblender: the Bayesian Light Source Separator (BLISS). BLISS builds on recent ideas in deep generative modeling and variational inference (Blei et al., 2017; Kingma and Welling, 2019) to quantify blending uncertainty. Propagating this uncertainty to downstream cosmological analyses has the potential to mitigate blending systematics.
@@
BLISS uses simulation-based inference (SBI) to approximate a posterior distribution of deblended galaxy properties (see Section 3). [^1] As explained in Section 3, the galaxy deblending component of BLISS is deterministic and uses an autoencoder architecture (Section 3.4). Despite this, BLISS can produce posterior distributions of galaxy properties that capture detection uncertainty using the procedure described in Section 4.6. BLISS extends the approach from StarNet (Liu et al., 2023), which uses SBI to detect and measure stars in crowded starfields, to process images containing both stars and galaxies. Algorithmically, inference involves three encoders and a decoder, as illustrated in Figure 1. After training these models, inference can be performed rapidly on a GPU, in just seconds for megapixel images. In this work, we train and evaluate BLISS on simulated single-band images that are representative of typical year-10 LSST wide-field coadd images used for cosmological analyses.
@@ 2 Simulated Images
We designed our dataset of simulated images to mimic year-10 LSST single-band coadds. The galaxies and stars in our image simulations are modeled using parametric light profiles. The source density is taken from previous LSST simulation work (Connolly et al., 2014; LSST-DESC et al., 2021), and we do not model clustering. All our images are simulated using the GalSim Python package (Rowe et al., 2015). [^2] https://github.com/GalSim-developers/GalSim
@@ 2.1 Survey Settings
Our studies target a single band, the $i$-band, at the LSST survey pixel scale of 0.2 arcseconds. We use the same parametric point spread function (PSF) model $\pi$ for all our galaxy images, which consists of an atmospheric Kolmogorov component and an optical obscured Airy pattern. We use a FWHM of $0.79$ for the atmospheric component, corresponding to the median zenith seeing for this filter (Ivezić et al., 2019a). We add Gaussian noise to our images, assuming the background-dominated limit, so that their SNR resembles that of LSST year-10 images produced by coaddition. The added noise corresponds to a sky level $b$, which is chosen to be the average expected sky level for LSST year 10. However, the sky level is not added directly to the images; all simulated images are background-subtracted. Our setup closely resembles the simulations used in previous blending work focused on LSST (Sanchez et al., 2021a). The exact survey parameters are sourced from the LSST module of SurveyCodex. [^3] SurveyCodex is a Python library designed to archive the parameters of various surveys with references. Many of its parameters were obtained from the WeakLensingDeblending package (Sanchez et al., 2021b), in some cases with corrections. This library can be found in the following URL: https://github.com/LSSTDESC/surveycodex
@@ 2.2 Galaxy Properties
The galaxy properties for our image simulations come from a galaxy catalog produced with CatSim, an early version of an end-to-end LSST simulation framework (Connolly et al., 2014). The galaxies produced by this framework cover the redshift range of $0<z<6$. They are simulated using a parametric model for the light profile of galaxies consisting of bulge and disk components. [^4] The disk component corresponds to a Sérsic profile with index $n=1$ (also known as an exponential profile), and the bulge component corresponds to a Sérsic profile with index $n=4$ (also known as a “de Vaucouleurs” profile) (Sérsic, 1963). The model parameters are the total flux, the relative flux of each component, and the sizes and shapes of each component. The two components share the same centroid and orientation. While CatSim also includes an AGN component, we omit it to simplify the modeling of galaxy morphology. The fluxes, shapes, and sizes of galaxies follow a realistic distribution designed to match LSST’s key observables, and are informed by previous SDSS observations as well as galaxy formation literature (Bruzual and Charlot, 2003; De Lucia et al., 2006; González et al., 2009).
For our study, we used the same one-square-degree subset of the CatSim catalog as in Sanchez et al. (2021a). This subset of galaxies comprises approximately 850850k galaxies with a limiting $r$-band magnitude of $28$. For our datasets, we impose a magnitude cut in the $i$-band of $27$, which corresponds to galaxies with a median SNR of approximately $3$. This yields approximately 576576k galaxies, which we split into three equal sets for training, validation, and testing.
@@ 2.3 Star Properties
Our datasets include stars to evaluate our deblender in the context of star-galaxy blends. The star distribution and fluxes are derived from the LSST DESC Data Challenge Two (DC2) simulation catalogs (LSST-DESC et al., 2021). Specifically, we use the same star catalog as in previous work on cosmic shear measurements in LSST-like simulations (Sheldon et al., 2023). We set our star density to the average over simulated fields, while rejecting fields with density higher than $100\,\text{arcmin}^{-2}$. This results in a star density of $15.5\,\text{arcmin}^{-2}$. We apply a lower limit magnitude cut of $20$ and an upper limit magnitude cut of $27$ in the $i$-band, resulting in approximately 1717k stars. The former cut excludes stars with very high SNRs of over $1000$. We model stars as a point source convolved by the same PSF $\pi$ used by galaxies. We do not simulate saturation or other kinds of image artifacts.
@@ 3 Method
BLISS estimates the posterior distribution of light source properties from a given astronomical image, following the procedure laid out in Figure 1. We start by decomposing the inference problem into subproblems involving overlapping image subregions, which we call “padded tiles” (Section 3.1). These padded tiles are passed into a set of three neural network encoders that are trained via variational inference (Ambrogioni et al., 2019) to return the parameters of distributions characterizing the light source properties in each tile (Section 3.2). First, the detection encoder returns posterior distributions on the number of objects per tile and the positions of objects’ centroids. Sampled centroids from these distributions are used to center the padded tiles (within a pixel, via cropping). Next, the classification encoder takes these centered padded tiles and returns a probability characterizing the type of each source (star or galaxy) (Section 3.3). Finally, the deblending encoder uses these centered padded tiles to return (deterministic) latent galaxy representations that can be used to obtain noiseless reconstructions of deblended galaxy images (Section 3.4). Once the three encoders have been trained, BLISS can process astronomical images of any size with the same PSF and resolution as the training images.
@@ 3.1 Tiling
BLISS, by construction, outputs a prediction for at most one source within a given tile. As a result, if the centroids of two or more sources are present in a single tile, BLISS will output an incomplete catalog for that tile. In particular, if at least one source in the tile is detectable, BLISS will output a set of properties corresponding to a single “hybrid” source. [^8] At training time, BLISS always targets the properties of the brightest source present in a tile. We partially mitigate this by using small tiles of size 5 pixels, which reduces the probability that more than one source centroid lies in any given tile. Given a source density of $175.5\,\rm{arcmin}^{-2}$, this probability is approximately $0.0012$. To be concrete, we expect that approximately $10\%$ of images in our dataset-blend dataset will have at least one tile with two source centroids. In the majority of these occurrences, however, one of the sources involved will be below the detection threshold. [^9] Realistic levels of clustering might significantly impact these assumptions. Clustering increases the number of close pairs of objects at a given number density (Euclid Collaboration et al., 2019; MacCrann et al., 2022). This would increase the proportion of objects that fall in the same tile, thus degrading BLISS performance in the manner mentioned above.
@@ 3.2 Variational inference
BLISS adopts a standard Bayesian model of astronomical images, whereby light source properties are latent random variables, and pixel intensities are observed random variables (cf. Brewer et al., 2013; Portillo et al., 2017; Regier et al., 2019; Buchanan et al., 2023; Stone et al., 2023). Appendix A provides details of our specific Bayesian model.
@@
In our approximation, the inferred tile catalog $\mathcal{Z}_{t}$ contains the light source properties $\{n_{t},\bm{\ell}_{t},b_{t},\bm{z}_{t}\}$, where $n_{t}\in\{0,1\}$ is the number of sources whose centroids lie in the tile, $\bm{\ell}_{t}\in[0,1]^{2}$ is the centroid position of the tile’s source in units of tile length, $b_{t}\in\{0,1\}$ is the classification of the source as a galaxy ($b_{t}=1$) or star ($b_{t}=0$), and $\bm{z}_{t}\in\mathbb{R}^{8}$ is a latent vector characterizing galaxy morphology and brightness (for more details, see Section 3.4).
@@
are the output of neural networks with weights $\bm{\phi}$ applied to a given padded tile $\mathcal{X}_{t}$. These weights are the variational parameters. We model locations using a bivariate Gaussian in the variational distribution. Samples from this distribution that lie beyond the bounds of a given tile are considered non-detections. [^10] When the true centroid is close to the tile boundary or highly uncertain, there is a significant probability that the posterior samples fall outside the tile. We choose to discard these samples instead of assigning this source to a separate tile due to our current simplifying assumption that tiles are completely independent. Future work will explore sharing information between neighboring tiles to improve the variational approximation. Additionally, we chose the variational distribution for the latent galaxy representation $\bm{z}_{t}$ to be a delta function. This implies that we only return point estimates for this quantity with no associated uncertainty.
Critically, our method is amortized (Kingma and Welling, 2013; Rezende et al., 2014): the variational parameters $\bm{\phi}$ are shared among tiles (i.e., $\bm{\phi}$ is not indexed by $t$); a single set of neural networks with weights $\bm{\phi}$ is used for any padded tile of any image. If our method were not amortized, as in more classical variational inference approaches (Blei et al., 2017), we would need to run an iterative optimization procedure for each tile, which would make our procedure highly inefficient. Amortization allows us to train a single set of encoders only once, which greatly speeds up inference for large datasets.